3. At 300 K, 0.045 mol of hydrogen and fluorine are placed in a closed 1.5 L container and allowed to reach equilibrium.

               H2(g) + F2(g) <---->  2HF(g)                     Keq = 4.3 at 300 K


What is the equilibrium concentration of HF(g)?  


1
Expert's answer
2021-01-15T06:30:35-0500

Initial concentration of H2 and F2 will be 0.045/1.5 = 0.03 M


"4.3 = \frac{(2x)^2}{(0.03-x)(0.03-x)} \\

\sqrt{4.3} = \frac{2x}{0.03-x} \\

2.073 = \frac{2x}{0.03-x} \\

0.06219 -2.077x = 2x \\

0.06219 = 4.077x \\

x = 0.015 \;M"

The equilibrium concentration of HF is "2 \times 0.015 = 0.03 \; M"

Answer: 0.03 M